Find the probability that the sum is as stated when a pair of dice is rolled. (Enter your answers as fractions.)
A. even and doubles
B. even or doubles

Answer by mananth(16947)

.
The above shows the 36 outcomes when two dice are rolled
Pick up the conditions stated and find probability.
P= Number of possible outcomes/ Total number of outcomes.
Choose from above
Doubles (2,2),(4,4),(6,6)
P = 3/36 = 1/12
Doubles or even
You locate there are 24 outcomes
P = 24/36 =2/3